One leg of an isosceles right triangle measures 12 feet. Find the exact length of the hypotenuse.
Answer by stanbon(75887)

If one leg is 12, the other length is 12.
Use Pythagoras to find the hypotenuse:
h^2 = 12^2 + 12^2
h^2 = 2*12^2
h = 12*sqrt(2)
==================
Cheers,
Stan H.